# Think Natively: Unlocking Multilingual Reasoning with Consistency-Enhanced Reinforcement Learning

### v2

@@ Abstract
Large Reasoning Models (LRMs) have achieved remarkable performance on complex reasoning tasks by adopting the “think-then-answer” paradigm, which enhances both accuracy and interpretability. However, current LRMs exhibit two critical limitations when processing non-English languages: (1) They often struggle to maintain input-output language consistency; (2) They generally perform poorly with wrong reasoning paths and lower answer accuracy compared to English. These limitations significantly degrade the user experience for non-English speakers and hinder the global deployment of LRMs. To address these limitations, we propose M-Thinker, which is trained by the GRPO algorithm that involves a Language Consistency (LC) reward and a novel Cross-lingual Thinking Alignment (CTA) reward. Specifically, the LC reward defines a strict constraint on the language consistency between the input, thought, and answer. Besides, the CTA reward compares the model’s non-English reasoning paths with its English reasoning path to transfer its own reasoning capability from English to non-English languages. Through an iterative RL procedure, our M-Thinker-1.5B/7B models not only achieve nearly 100% language consistency and superior performance on two multilingual benchmarks (MMATH and PolyMath), but also exhibit excellent generalization on out-of-domain languages.
@@ 1 Introduction
However, current LRMs generally suffer from two major issues under multilingual scenarios. First, they often struggle to maintain input-output language consistency Wang et al. (2025d); Tam et al. (2025), i.e., they frequently default to thinking and answering in English (or other unintended languages) rather than the input language (please refer to Figure 1). Second, they present inferior performance for low-resource languages compared to English Luo et al. (2025); Wang et al. (2025d). These issues significantly reduce the readability of the reasoning process and degrade the user experience of LRMs in multilingual environments. To mitigate these issues, current solutions include language control instructions Tam et al. (2025), supervised fine-tuning (SFT) with specific language data Luo et al. (2025), and GRPO Shao et al. (2024) with a soft language reward Park et al. (2025); Mistral-AI (2025); Hwang et al. (2025). However, these solutions still face notable limitations: Prompt-based methods struggle to enforce output language consistency with the input; SFT generally entails a trade-off between answer accuracy and language consistency; Soft consistency rewards in GRPO can only impose weak constraints on maintaining language consistency. Therefore, there still remains a clear need for a solution to effectively enhance both language consistency and multilingual reasoning capability of LRMs.
@@
Experimental results on two publicly-used multilingual benchmarks (MMATH and PolyMath) show that our M-Thinker-1.5B/7B models not only achieve nearly 100% language consistency and substantial performance improvement, but also demonstrate remarkable generalization on out-of-domain languages. In summary, the major contributions of this paper are as follows [^1] https://github.com/XZhang00/M-Thinker:
@@
  • Experimental results of our M-Thinker-1.5B/7B models on MMATH and PolyMath benchmarks demonstrate superior performance on both language consistency and answer accuracy for multiple languages.
@@ 4.1 Experimental Setups
We select two commonly-used models with different sizes as our backbones: DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025). The two models exhibit imbalanced reasoning performance in different languages, showing better ability in English and Chinese compared to other languages. Based on the imbalanced ability of the two models and the included languages of the MMATH Luo et al. (2025) benchmark, we select Japanese (ja), Korean (ko), French (fr), Portuguese (pt), and Thai (th) as the training (in-domain, ID) languages and English (en), Spanish (es), Arabic (ar), Vietnamese (vi), and Chinese (zh) as out-of-domain (OOD) languages to observe the generalization [^3] Since the original model performs well on en and zh, we actually want to observe the catastrophic forgetting phenomenon for en and zh. To simplify writing, we refer to it as generalization here. of each method. The details for each language are introduced in Table 6 of Appendix B.1.
@@
We set the iterations for RL training $I$ is 2. The detailed training settings of cold-start SFT and iterative RL training, and evaluation details are listed in Appendix B.3.
4.2 Baselines
Prompt-Control.
@@ 4.2 Baselines
DIT (No Training) 1.96 0.02 7.24 25.78 0.66 7.13 91.07 15.91 1.23 13.42 64.25 37.17 22.15
QRT (No Training) 10.69 0.22 27.01 45.98 5.96 17.97 92.37 45.89 5.21 37.14 65.37 49.19 33.58
Metric: Accuracy (Acc, %)
DIT (No Training) 37.06 34.18 39.53 34.00 29.29 34.81 45.97 43.65 32.50 33.09 39.95 39.03 36.92
QRT (No Training) 30.06 34.93 36.33 29.58 25.60 31.30 46.28 25.74 29.49 22.82 40.04 32.87 32.09
Metric: Language Consistency & Accuracy (LC&Acc, %)
DIT (No Training) 0.32 0.00 4.18 14.58 0.12 3.84 45.93 11.69 0.40 4.82 33.76 19.32 11.58
QRT (No Training) 4.94 0.14 13.10 23.52 0.52 8.45 46.22 23.75 1.39 13.91 35.42 24.14 16.29
DIT.
QRT.
Question-Restatement Thinking Luo et al. (2025) restates the question in the target language at the beginning of the thinking process, which encourages the model to generate the thinking content in the target language. The restatement instructions for each language are listed in Figure 4 of Appendix B.4.
Cold-Start SFT.
Naive-RL.
We equip the GRPO algorithm only with the accuracy reward to conduct the RL training. The training dataset is the same as our first training iteration (Iter-1).
SLC-RL.
We equip the GRPO algorithm with the accuracy reward and a soft language consistency reward Mistral-AI (2025) to conduct the RL training, i.e., $R(o)=R_{\text{acc}}(o)+R_{\text{slc}}(o)$. When the answer is correct $R_{\text{acc}}(o)=0.9$, and when the language is consistent with the input language $R_{\text{slc}}(o)=0.1$, otherwise, $R_{\text{acc}}(o)=R_{\text{slc}}(o)=0$. The training dataset is the same as our first training iteration (Iter-1).
4.3 Main Results
Performance of our M-Thinker.
We report the evaluation results on MMATH of the DeepSeek-R1-Distill-Qwen-1.5B/7B backbones in Table 1 and Table 2. The results demonstrate that our M-Thinker-1.5B/7B achieves excellent improvement on LC, Acc, and the combined metric (LC&Acc). On the main evaluation metric (LC&Acc), our M-Thinker-1.5B/7B (Iter-1) drastically outperforms all baselines, which highlights the effectiveness of our designed rewards in simultaneously optimizing for correctness and language fidelity. Surprisingly, the performance on MMATH of our M-Thinker-7B (Iter-1/2) even outperforms DeepSeek-R1-0528 on LC&Acc (particularly in fr/pt/th/es/vi, as shown in Table 2), indicating the powerful multilingual reasoning ability of our M-Thinker-7B. Furthermore, our M-Thinker-1.5B/7B (Iter-2) achieves further improvement than Iter-1, which proves that our iterative training procedure can progressively enhance the model’s capabilities. And the performance on LC&Acc of our M-Thinker-1.5B/7B (Iter-2) has surpassed the performance on Acc of the backbones DeepSeek-R1-Distill-Qwen-1.5B/7B, which means that utilizing the input language to respond can exceed the performance of responding in English or other default languages. This superior performance indicates that our method almost overcomes the trade-off between language consistency and answer accuracy, improving both language consistency and answer correctness to achieve powerful multilingual reasoning ability.
In-Domain Languages Out-of-Domain Languages
Methods ja ko fr pt th ID-AVG en es ar vi zh OOD-AVG ALL-AVG
Metric: Language Consistency (LC, %)
DeepSeek-R1-0528 70.44 65.17 44.72 43.16 23.37 49.37 72.56 37.95 64.03 13.60 69.22 51.47 50.42
@@ 4.3 Main Results
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 98.69 99.82 99.56 99.32 99.52 99.38 99.98 99.42 82.06 100.00 90.15 94.32 96.85
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 59.95 56.06 65.61 67.24 60.24 61.82 71.86 64.89 62.36 60.53 67.92 65.51 63.67
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 59.87 55.89 65.59 66.77 60.18 61.66 71.84 64.73 56.30 60.53 60.68 62.81 62.24
Table 2: The LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the DeepSeek-R1-Distill-Qwen-7B backbone. Other symbols have the same meaning as in Table 1.
Performance of baselines.
No training baselines have a minor improvement on LC&Acc, and QRT outperforms Prompt-Control and DIT. The performance of these prompt-based methods heavily depends on the original instruction-following ability of backbones, i.e., the larger improvement on 7B than 1.5B. Additionally, the improvement on LC and the decrease on Acc also reflect the trade-off between the language consistency and answer accuracy. Naive-RL (GRPO only with the accuracy reward) shows the best results on Acc but the lowest LC (0.0) since the responses generated in English can obtain a higher reward score during RL training, so that the trained model is most likely to answer in English, which is contrary to the goal of a multilingual reasoning model. Although SLC-RL is trained with a soft language consistency reward, the models still struggle to maintain language consistency, particularly for the 1.5B backbone.
@@
Refer to the “OOD-avg”, our M-Thinker-1.5B/7B also significantly surpasses other baselines, which indicates that the reasoning patterns learned through our rewards and training procedure are not confined to the training languages but are successfully transferred to unseen languages. The evaluation results on PolyMath (as shown in Table 7 in Appendix C) also present similar trends, which further prove the superiority of our method.
In summary, these results demonstrate that M-Thinker-1.5B/7B effectively improves both the language consistency and answer accuracy in multilingual reasoning scenarios.
Judge Models ID-AVG OOD-AVG ALL-AVG
w/o $R_{\text{cta}}$ 31.68 37.18 34.43
DeepSeek-V3-0324 35.39 38.37 36.88
Qwen2.5-7B-Instruct 31.69 34.13 32.91
Table 4: The LC&Acc results of different judge models for $R_{\text{cta}}$ based on our M-Thinker-1.5B (Iter-1).
@@ 5.1 Ablation Study
We conduct an ablation study to verify the effectiveness of our designed reward functions and involved training strategies. The ablation results listed in Table 3 show that the LC&Acc performance degrades in both ID and OOD languages without $R_{\text{cta}}$. For the setting “w/o $R_{\text{lc}}$”, although the Acc improves over M-Thinker-1.5B, the model responds to all questions in English, resulting in the lowest language consistency. “w/o ($R_{\text{cta}}$ & $R_{\text{lc}}$)” present the lowest performance. These results prove the effectiveness of our designed reward functions. Additionally, “w/o Cold-Start SFT” and “w/o Rejection Sampling” also have a performance decline, which demonstrates the necessity of these strategies. Furthermore, directly using English responses from the Light-R1-SFT dataset for $R_{\text{cta}}$ also underperforms our M-Thinker (using generated English responses from the model itself), since the latter may have a smaller gap between the abilities of non-English languages and English. Detailed results of each setting are listed in Table 8 of Appendix D.
5.2 Different Judge Models for $R_{\text{cta}}$
In this section, we report the different performance of utilizing different judge models to calculate $R_{\text{cta}}$ in Table 4. We find that prompting Qwen2.5-7B-Instruct to judge the alignment ratio decreases the overall performance (32.91%) compared to using DeepSeek-V3-0324 (36.88%), and even underperforms the setting without $R_{\text{cta}}$ (34.43%). These results demonstrate that Qwen2.5-7B-Instruct (with the smaller size) provides noise rewards when judging the alignment ratio, resulting in performance degradation than “w/o $R_{\text{cta}}$”. By contrast, DeepSeek-V3-0324 can provide effective rewards, achieving the best performance.

### v3

@@ Abstract
Large Reasoning Models (LRMs) have achieved remarkable performance on complex reasoning tasks by adopting the “think-then-answer” paradigm, which enhances both accuracy and interpretability. However, current LRMs exhibit two critical limitations when processing non-English languages: (1) They often struggle to maintain input-output language consistency; (2) They generally perform poorly with wrong reasoning paths and lower answer accuracy compared to English. These limitations significantly compromise the interpretability of reasoning processes and degrade the user experience for non-English speakers, hindering the global deployment of LRMs. To address these limitations, we propose M-Thinker, which is trained by the GRPO algorithm that involves a Language Consistency (LC) reward and a novel Cross-lingual Thinking Alignment (CTA) reward. Specifically, the LC reward defines a strict constraint on the language consistency between the input, thought, and answer. Besides, the CTA reward compares the model’s non-English reasoning paths with its English reasoning path to transfer its own reasoning capability from English to non-English languages. Through an iterative RL procedure, our M-Thinker-1.5B/4B/7B models not only achieve nearly 100% language consistency and superior performance on two multilingual benchmarks (MMATH and PolyMath), but also exhibit excellent generalization on out-of-domain languages.
@@ 1 Introduction
However, current LRMs generally suffer from two major issues under multilingual scenarios. First, they often suffer from input-output language inconsistency Wang et al. (2025d); Tam et al. (2025), i.e., they frequently default to thinking and answering in English (or other unintended languages) rather than the input language (please refer to Figure 1). Second, they present inferior performance for other languages compared to English Luo et al. (2025); Wang et al. (2025d). These issues significantly reduce the readability and explainability of the reasoning processes and degrade the user experience of LRMs in multilingual environments. To mitigate these issues, current solutions include language control instructions Tam et al. (2025), supervised fine-tuning (SFT) with specific language data Luo et al. (2025), and GRPO Shao et al. (2024) with a soft language reward Park et al. (2025); Mistral-AI (2025); Hwang et al. (2025). However, these solutions still face notable limitations: Prompt-based methods struggle to enforce output language consistency with the input; SFT generally entails a trade-off between answer accuracy and language consistency; Soft consistency rewards in GRPO can only impose weak constraints on maintaining language consistency. Therefore, there still remains a clear need for a solution to effectively enhance both language consistency and multilingual reasoning capability of LRMs.
@@
Experimental results on two publicly-used multilingual benchmarks (MMATH and PolyMath) show that our M-Thinker-1.5B/4B/7B models not only achieve nearly 100% language consistency and substantial performance improvement, but also demonstrate remarkable generalization on out-of-domain languages. In summary, the major contributions of this paper are as follows [^1] https://github.com/XZhang00/M-Thinker:
@@
  • Experimental results of our M-Thinker-1.5B/4B/7B models on MMATH and PolyMath benchmarks demonstrate superior performance on both language consistency and answer accuracy for multiple languages.
@@ 4.1 Experimental Setups
We select three commonly-used reasoning models with different sizes as our backbones: DeepSeek-R1-Distill-Qwen-1.5/7B DeepSeek-AI (2025) and Qwen3-4B-Thinking-2507 Yang et al. (2025a). The three models exhibit imbalanced reasoning performance in different languages, showing better ability in English compared to other languages. Based on the imbalanced ability and the included languages of the MMATH Luo et al. (2025) benchmark, we select Japanese (ja), Korean (ko), French (fr), Portuguese (pt), and Thai (th) as the training (in-domain, ID) languages and English (en), Spanish (es), Arabic (ar), Vietnamese (vi), and Chinese (zh) as out-of-domain (OOD) languages to observe the generalization [^3] Since the original model performs well on en, we actually want to observe the catastrophic forgetting phenomenon for en. To simplify writing, we refer to it as generalization here. of each method. The details for each language are introduced in Table 7 of Appendix B.1.
@@
We set the iterations for RL training $I$ is 2. The detailed training settings of cold-start SFT and iterative RL training, and generation configs are listed in Appendix B.3.
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 97.86 99.37 99.50 99.05 95.68 98.29 98.00 99.44 75.12 100.00 90.97 92.70 95.50
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 58.23 60.56 68.58 66.99 63.98 63.66 71.75 68.34 63.00 61.72 67.25 66.41 65.04
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 57.50 60.26 68.52 66.87 63.44 63.32 71.71 68.22 53.70 61.72 60.58 63.18 63.25
Table 1: The LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the DeepSeek-R1-Distill-Qwen-7B backbone. “ID-avg/OOD-avg” is the average result of five In-Domain/Out-of-Domain languages and “ALL-AVG” is the average result of all ten languages. The result in bold means the best result, and the underlined result means the second-best result in each setting. “Iter-1/2” means the training iteration 1/2.
4.2 Baselines
Prompt-Control.
DIT.
QRT.
Question-Restatement Thinking Luo et al. (2025) restates the question in the target language at the beginning of the thinking process, which encourages the model to think in the target language. The restatement instructions for each language are listed in Figure 4 of Appendix B.4.
Cold-Start SFT.
Naive-RL.
We equip the GRPO algorithm only with the accuracy reward to conduct the RL training based on the same cold-started SFT model. The training dataset is the same as our first training iteration (Iter-1).
SLC-RL.
We equip the GRPO algorithm with the accuracy reward and a soft language consistency reward Mistral-AI (2025) to conduct the RL training, i.e., $R(o)=R_{\text{format}}(o)*(R_{\text{acc}}(o)+R_{\text{slc}}(o))$. When the format is correct: $R_{\text{format}}(o)=1$, when the answer is correct: $R_{\text{acc}}(o)=0.9$, and when the language is consistent with the input language: $R_{\text{slc}}(o)=0.1$, otherwise, $R_{\text{format}}(o)=R_{\text{acc}}(o)=R_{\text{slc}}(o)=0$. The initial policy model (after cold-start SFT) and training dataset is the same as our first training iteration (Iter-1).
@@ 4.2 Baselines
DeepSeek-R1-Distill-Qwen-1.5B 5.98 36.11 21.04 34.81 39.83 37.32 3.87 19.19 11.53
Prompt-Control (No Training) 12.64 48.52 30.58 31.95 33.55 32.75 5.65 22.32 13.99
DIT (No Training) 22.48 43.30 32.89 23.10 28.50 25.80 11.56 23.68 17.62
QRT (No Training) 23.47 45.46 34.46 19.26 28.11 23.69 11.68 23.87 17.78
Cold-Start SFT 23.73 47.75 35.74 19.18 27.79 23.49 7.39 21.73 14.56
Naive-RL 0.00 30.97 15.48 49.99 50.08 50.04 0.00 16.16 8.08
SLC-RL 0.00 37.16 18.58 46.80 49.16 47.98 0.00 19.47 9.74
M-Thinker-1.5B $\Rightarrow$ Iter-2 (Ours) 99.49 79.51 89.50 42.72 46.53 44.62 42.47 38.83 40.65
Table 2: The LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the DeepSeek-R1-Distill-Qwen-1.5B backbone. The detailed results for each language are list in Table 9 of Appendix C.2.
LC Acc LC&Acc
Settings ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG
@@
4.3 Main Results
Performance of our M-Thinker.
We report the evaluation results on MMATH of the three backbones in Table 1, Table 2, and Table 8 (Appendix C.1). The results demonstrate that our M-Thinker-1.5B/4B/7B achieves excellent improvement on LC, Acc, and the combined metric (LC&Acc). On the main evaluation metric (LC&Acc), our M-Thinker-1.5B/4B/7B (Iter-1) drastically outperforms all baselines, which highlights the effectiveness of our designed rewards in simultaneously optimizing for correctness and language fidelity. Furthermore, our M-Thinker-1.5B/7B (Iter-2) achieves further improvement than Iter-1, which proves that our iterative training procedure can progressively enhance the model’s capabilities. And the performance on LC&Acc of our M-Thinker-1.5B/7B (Iter-2) has surpassed the performance on Acc of the backbones DeepSeek-R1-Distill-Qwen-1.5B/7B, which means that responding in the input language can exceed the performance of responding in English or other default languages. This superior performance indicates that our method mitigates the trade-off between language consistency and answer accuracy, achieving powerful multilingual reasoning ability.
Performance of baselines.
No training baselines have a minor improvement on LC&Acc, and QRT outperforms DIT and Prompt-Control. The performance of these prompt-based methods heavily depends on the original instruction-following ability of backbones, i.e., the larger improvement on 7B than 1.5B. Additionally, the improvement on LC and the decrease on Acc also reflect the trade-off between the language consistency and answer accuracy. Naive-RL (GRPO only with the accuracy reward) shows the best results on Acc but the lowest LC (0.0) since the responses generated in English can obtain a higher reward score during RL training, so that the trained model is most likely to think and answer in English, which is contrary to the goal of a multilingual reasoning model. Although SLC-RL is trained with a soft language consistency reward, the models still struggle to maintain language consistency, particularly for the 1.5B backbone. By contrast, our method with the strict LC reward can promote the input-output language consistency while having no degradation [^7] More detailed analyses about hard/soft LC reward are listed in Appendix C.3. on Acc compared to SLC-RL [^8] We also conduct SLC-RL with the same reward magnitude as ours and present it in Table 12 of Appendix C.4. (58.87 vs. 58.59).
@@ 4.3 Main Results
Refer to the “OOD-avg”, our M-Thinker also significantly surpasses other baselines, which indicates that the reasoning patterns learned through our rewards and training procedure are not confined to the training languages but are successfully transferred to unseen languages. The evaluation results on PolyMath (as shown in Table 13 in Appendix C.5) also present similar trends, which further prove the superiority of our method.
In summary, these results demonstrate that our M-Thinker effectively improves both the language consistency and answer accuracy in multilingual reasoning scenarios.
@@ 5.1 Ablation Study
We conduct an ablation study to verify the effectiveness of our designed reward functions and involved training strategies. The ablation results listed in Table 3 show that the LC&Acc performance degrades in both ID and OOD languages without $R_{\text{cta}}$. For the setting “w/o $R_{\text{lc}}$”, although the Acc improves over M-Thinker-1.5B, the model responds to all questions in English, resulting in the lowest language consistency. “w/o ($R_{\text{cta}}$ & $R_{\text{lc}}$)” present the lowest performance. These results prove the effectiveness of our designed reward functions. Additionally, “w/o Cold-Start SFT” and “w/o Rejection Sampling” also have a performance decline, which demonstrates the necessity of these strategies. Furthermore, directly using English responses from the Light-R1-SFT dataset (which is generated by DeepSeek-R1) for $R_{\text{cta}}$ also underperforms our M-Thinker (using generated English responses from the model itself), since the latter may have a smaller gap between the abilities of non-English languages and English. Detailed results of each ablation setting are listed in Table 14 of Appendix D.1.
5.2 Effects of Different Judge Models for $R_{\text{cta}}$
In this section, we analyze the effects of different judge models for calculating $R_{\text{cta}}$ on performance and report the results in Table 4.
Findings 1: Frontier small LLMs can also provide reliable rewards. Beyond DeepSeek-V3, we also test two smaller models as the judge model, i.e., Qwen3-30B-A3B and Qwen3-4B. Although smaller, these two frontier models still deliver notable performance gains while being more cost-efficient than DeepSeek. Besides, we also try another small model, Qwen2.5-7B-Instruct, that is relatively outdated compared to other models. We find that it decreases the overall performance (32.91%) compared to “w/o $R_{\rm cta}$” due to the limited multilingual capability, demonstrating that the multilingual capability of the judge model is crucial for the effectiveness of the $R_{\text{cta}}$ reward.
Findings 2: $R_{\text{cta}}$ achieves cross-lingual transfer of the reasoning capability from English to other languages. As shown in Table 4, we also find that $R_{\text{cta}}$ significantly brings the accuracy gap between English and other languages with multiple judge models according to the “GAP” values. This further proves the effectiveness of $R_{\text{cta}}$ for bridging the multilingual reasoning gap in existing models.
[…]
w/o $R_{\text{cta}}$ 31.68 37.18 34.43 13.47
DeepSeek-V3-0324 35.39 38.37 36.88 9.24
Qwen3-4B-Instruct-2507 32.48 38.51 35.49 12.56
Qwen3-30B-A3B-Instruct-2507 33.12 37.08 35.10 11.75
Qwen2.5-7B-Instruct 31.69 34.13 32.91 13.65
Table 4: The LC&Acc results of different judge models for $R_{\text{cta}}$ based on our M-Thinker-1.5B (Iter-1). “GAP” denotes the accuracy gap between English and other languages.
@@ 5.3 Generalization Study
5.4 Human Evaluation
To empirically validate the reliability of our utilized judge model (DeepSeek-V3-0324) for $R_{\text{cta}}$, we conduct a human evaluation on a randomly sampled subset of the training data (30 samples per language). We collaborate with a professional data annotation service to recruit linguistic experts proficient in the target languages (ja/ko/fr/pt/th). We request that they return the alignment ratios, and the detailed annotation guidelines are introduced in Appendix D.3. We then calculate the Pearson correlation between these human-annotated ratios and the judge model’s CTA scores. The results in Table 6 demonstrate a positive Pearson correlation coefficient across different languages, confirming that the judge model aligns well with human judgment and proving the reliability of the CTA reward.
Language from Human from LLM Pearson Correlation
ja 0.72 0.69 0.87
ko 0.68 0.64 0.86
fr 0.85 0.81 0.93
pt 0.82 0.78 0.91
th 0.63 0.57 0.83
AVG 0.74 0.70 0.88
Table 6: The Pearson correlation between human-annotated ratios and the judge model’s CTA scores (DeepSeek-V3-0324).